OpenStudy (brittneygarcia5):

how do you determine algebraically if an equation is one-to-one? For example: h(x)=4/(x^2)

OpenStudy (mathmate):

If f(x1)=f(x2) where \(x1\ne x2\) , then f(x) is not one to one. Similarly, if f(x1) = y1 and f(x2)=y2, where x1=x2, then f(x) is not a function.

OpenStudy (brittneygarcia5):

so, you plug in different numbers, and if two match up, then it is not a one-to-one function, right? @mathmate

OpenStudy (mathmate):

Yeah, but do not plug in numbers at random! lol Examine the function, and see if you can find distinct numbers x1 and x2 such that f(x1)=f(x2).

OpenStudy (brittneygarcia5):

so, how do you know which numbers to plug in?

OpenStudy (brittneygarcia5):

@mathmate

OpenStudy (mathmate):

Hint: Are you aware that \((-5)^2=(5)^2\)

OpenStudy (brittneygarcia5):

oh yeah duh, lol. so, look for patterns in the equation?

OpenStudy (mathmate):

Exactly. Also, your post requests algebraic way of proving it one-to-one, but nothing stops you from using the horizontal line test as a check.

OpenStudy (brittneygarcia5):

Thank you so much! @mathmate

OpenStudy (mathmate):

You're welcome! :)